what is the inverse of f(x) = 5x-10?
Answer by edjones(8007)

f(x) = 5x-10
y=5x-10
x=5y-10 the key step, switch x and y. Then solve for y.
5y=x+10
y=x/5 +2
f^(-1)(x)=x/5 +2
.
Ed